# A kinetic model of solar wind acceleration driven by ambipolar electric potential and velocity-space diffusion

### v2

@@ Abstract
Context. Parker Solar Probe (PSP) observations revealed that most of the solar wind acceleration occurs very close to the Sun. This acceleration is partly due to the global electric potential originating from the mass disparity between electrons and protons, coupled with the constraints of charge quasi-neutrality and zero-current conditions in the solar wind plasma. However, the exact mechanism that accounts for the remaining acceleration is still not identified.
Aims. We aim to provide a framework incorporating the electric field-driven component of the acceleration while also introducing an additional acceleration mechanism via velocity-space diffusion of the particles. This will help us determine the extent of extra acceleration, beyond the electric field-driven component, required to fully reproduce the acceleration of the solar wind in theoretical models.
Methods. We modified an existing kinetic exospheric model to account for the unexplained solar wind acceleration by including velocity-space diffusion, capturing the effect of collisions and wave-particle interactions within the exospheric approach. We compared the electric field derived from the sunward deficit of velocity distribution functions observed by PSP between 13.3 and 50 solar radii ($R_{s}$) with the electric field found self-consistently by the kinetic exospheric model.
Results. The effect of velocity-space diffusion is found to reduce the temperature anisotropy and impact the solar wind acceleration while maintaining the electric potential unchanged. The approach described in this work allows for adjusting the diffusion to effectively reduce or increase the solar wind acceleration. It is found that, even without diffusion, the model is able to reproduce the anticorrelation between the electric potential and the solar wind terminal velocity found by PSP. This suggests that the electric potential might still be of major importance in explaining the solar wind acceleration.
@@ 1 Introduction
The electrons and protons that are blown away from the Sun through interplanetary space are the main constituents of a continuous outflow of particles referred to as the solar wind. The exact mechanism that allows the solar wind to reach observed velocities averaging between 300 and 800 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ remains one of the most significant open questions in space physics (e.g., Rouillard et al. 2021). Since the first in situ solar wind observations in 1959, neither magnetohydrodynamic (MHD) nor kinetic models have been able to self-consistently explain the solar wind acceleration. In particular, MHD models are not able to capture the origin of the fast solar wind with realistic coronal temperatures and often require additional heating mechanisms (Parker 1965; Leer et al. 1982; Hansteen & Velli 2012). Although some kinetic approaches can provide an explanation for the presence of peculiar observed velocity distribution functions (VDFs) of solar wind particles and propose potential mechanisms of solar wind acceleration (e.g., Hollweg & Isenberg 2002; Vocks et al. 2005; Seough et al. 2015), they generally lack any global electric field-driven component of the acceleration. In this paper, we investigate this acceleration by improving a specific class of kinetic models known as exospheric models, which are fundamentally based on considering the electric field-driven component as the primary acceleration mechanism. A comprehensive review of these models has been provided by Fahr & Shizgal (1983).
@@
The first exospheric models were developed for planetary exospheres and were later adapted for the solar wind (Chamberlain 1960). These models describe the radial evolution of the particle VDFs by neglecting collisions between particles. The first model of the solar wind that included a consistent “ambipolar” electric field was developed by Lemaire & Scherer (1971) in a collisionless regime considered above the solar surface, at a given altitude called the “exobase”, which was chosen at 6 solar radii ($R_{s}$). This electric field is called ambipolar in the sense that it satisfies the quasi-neutrality and zero-current conditions between protons and electrons, observed to be preserved in the solar wind (e.g., Feldman et al. 1975). The challenging aspect of these models was to determine the radial evolution of this ambipolar electric field self-consistently with the exospheric approach, especially in the case of non-monotonic proton potential (typically found when assuming an exobase below 6 $R_{s}$, which better describes the solar wind case), as noted by Jockers (1970). Inspired by the latter, a generalized method to approximate the solution for this self-consistent electric potential was later developed by Lamy et al. (2003).
@@
The use of Kappa (also known as Lorentzian) VDFs for electrons, in contrast to Maxwellian VDFs for protons, constitutes one of the major breakthroughs that was first introduced in an exospheric model by Pierrard & Lemaire (1996) and applied to the solar wind by Maksimovic et al. (1997a). This approach was motivated by the need to account for the so-called suprathermal electrons observed extensively in the solar wind (e.g., Lin 1980; Pilipp et al. 1987; Pierrard et al. 2016; Abraham et al. 2022; Salem et al. 2023). Such suprathermal electron populations are better described by a Kappa distribution function rather than a single Maxwellian distribution, as shown, for instance, by Maksimovic et al. (1997b) and more recently by Zheng et al. (2024). These electrons are high-energy particles that do not follow the classical single Maxwellian distribution fitted to the lower-energy part (called the core) of the observed distribution. These suprathermal distributions are typical for the halo, which extends in all directions, and the electron strahl, a more concentrated component observed in the anti-sunward direction (e.g., Feldman et al. 1975; Pilipp et al. 1987; Pierrard et al. 2001b). The mathematical description of Kappa distributions required for exospheric models was first introduced by Pierrard & Lemaire (1996) and generalized for non-monotonic potentials by Lamy et al. (2003). Within the scope of this model, using a Kappa distribution for electrons increases the number of electrons with high speeds in the anti-sunward direction, which effectively increases the electric potential difference between any two radial distances (see Pierrard & Lazar 2010 for a review). Therefore, this allows more protons to reach the maximum of the non-monotonic proton potential (usually within 1 to 5 $R_{s}$), which in turn accelerates the wind to realistically high bulk speeds without requiring unrealistically high temperatures. Following the same approach, Zouganelis et al. (2004) showed that similar results could also be obtained with a sum of two Maxwellians or a sum of Maxwellian and Kappa distributions (for core and strahl), which would not limit the significance of the strahl to the power law factor $\kappa$.
@@
Given the simplicity and effectiveness of the approach of Lamy et al. (2003) in assigning at least part of the unexplained acceleration to the presence of suprathermal electrons, the model was later extended to minor ions (low abundance ions compared to protons) by Pierrard & Pieters (2014). More recently, an improved version of this exospheric model, incorporating regularized Kappa distribution functions (allowing for even lower $\kappa$ values and therefore a higher fraction of suprathermal particles), was compared with PSP and Solar Orbiter (SolO) observations by Pierrard et al. (2023). The results demonstrated that this model could reproduce the radial evolution of the solar wind bulk velocity but failed to accurately reproduce the observed anisotropy in the solar wind. This discrepancy arises from the intrinsic conservation of energy and magnetic moment, which strictly constrain modeled VDFs in velocity space.
@@
It has already been noted on multiple occasions that the idealized view provided by the exospheric model is not completely accurate. In particular, collisions and plasma waves are expected to be the mechanisms responsible for the scattering of the electron VDF in the solar wind (Pierrard et al. 2001c; Salem et al. 2003; Pierrard et al. 2011; Boldyrev et al. 2020; Micera et al. 2020; Verscharen et al. 2022; Yoon et al. 2024). These effects inevitably populate regions of velocity space that would otherwise be inaccessible and are believed to be the main mechanisms responsible for the presence of the so-called halo population of higher-energy electrons that are not preferentially anti-sunward directed, unlike the strahl (Feldman et al. 1975; Pilipp et al. 1987; Maksimovic et al. 1997b; Štverák et al. 2008, 2009). Therefore, we propose a way to simulate the effect of particle scattering on the moments of the electron VDF that are modeled by introducing a new, “diffused” population of particles.
@@
Our approach aims to improve the existing kinetic exospheric model by first properly describing the exospheric framework applied to the solar wind. Therefore, Section 2 provides a review of an existing kinetic exospheric model. Section 3 describes how we incorporate the velocity-space diffusion within the reviewed kinetic exospheric approach. In particular, we investigate how diffused electrons affect the solar wind temperature, bulk velocity, and density in Section 4. The relevance of the kinetic exospheric model is also addressed in Section 5, as it effectively explains one of the components responsible for solar wind acceleration through its self-consistent electric field, derived from the quasi-neutrality and zero-current conditions. Finally, the conclusions and future perspectives are provided in Section 6.
@@ 2 Review of the solar wind kinetic exospheric model
The kinetic exospheric model is based on the assumption that the solar wind becomes collisionless above a radial distance $r_{0}$, called the “exobase”. This is generally defined by considering that the ratio of the plasma density scale height over the particle mean free path (called the Knudsen number, Kn) must be equal to one for this assumption to be valid. More precisely, protons and electrons in the solar wind have different reference altitudes at which their respective Kn reaches unity, as electrons typically have a shorter mean free path than protons (Maksimovic et al. 1997a). However, for simplicity, we will not make this distinction here, considering a simple, sharp transition between collision-dominated and collisionless regions at $r_{0}=1.2\leavevmode\nobreak\ R_{s}$. This exobase level is arbitrarily chosen to emphasize the model’s physics while remaining generally close to the expected position of the actual exobase above the solar surface (Pierrard et al. 2023). We begin our mathematical formulation of the model from the general kinetic equation.
@@
is the force applied to each particle with ${g}$ being the gravitational acceleration, $q_{i}$ the charge of the particle, ${E}$ the electric field, and ${B}$ the magnetic field. The right-hand side of Equation (1) corresponds to particle–particle interactions. Considering the collisionless assumption in the exosphere (above the exobase), we can reduce this equation to the Vlasov equation,
@@
The Liouville theorem asserts that the VDF is constant along the trajectories of the system. This means that any function $f(c_{1},c_{2},...,c_{n})$, where $c_{1},c_{2},...,c_{n}$ are constants of the motion through external forces ${F}$, is a solution to Equation (3). In the case of the solar wind, the constants of the motion for each particle are the total (kinetic plus potential) energy,
@@
where $N_{p}$ is a factor of normalization that allows to retrieve the specified proton density at $r_{0}$ by integrating the proton VDF in velocity space, $w_{p}=\sqrt{2k_{b}T_{0p}/m_{p}}$ is the proton thermal velocity with $k_{b}$ being the Boltzmann constant, $T_{0p}$ being the proton temperature at $r_{0}$, $m_{p}$ being the proton mass, and $\mathcal{E}_{p}/(k_{b}T_{0p})$ is the normalized proton energy so that
@@
Here, $\kappa$ is the parameter associated with a power law decreasing Kappa VDF, instead of an exponentially decreasing Maxwellian VDF (i.e., for $\kappa\rightarrow\infty$ the Maxwellian distribution is retrieved); $w_{e}=\sqrt{(2\kappa-3)k_{b}T_{0e}/(\kappa m_{e})}$ is the electron thermal velocity, $T_{0e}$ is the electron temperature at $r_{0}$, $m_{e}$ is the electron mass, $\Gamma$ is the Gamma function, $N_{e}$ is a factor of normalization (defined in Section 3) that allows for retrieving the specified electron density at $r_{0}$ by integrating the electron VDF in velocity space, and $\mathcal{E}_{e}/[(\kappa-3/2)k_{b}T_{0e}]$ is the normalized electron energy so that
@@
being the normalized total potential for electrons. Note that $f_{p}$ and $f_{e}$ are functions of $\mathcal{E}$ but also $\mathcal{M}$ since the magnetic moment will have an influence on the allowed perpendicular velocities of particles depending on their origin.
@@
These equations are given in a general form for electrons and protons, as noted by the species index $i$ which will be omitted, for convenience, in the remainder of the description. If we assume that the reference altitude $r_{0}$ is exactly the point where the solar wind becomes collisionless and that below $r_{0}$ the particles are dominated by collisions (in such a way that there is no preferential direction for the particle motion), then at $r_{0}$ all particles should be constrained to $v_{\parallel}(r_{0})\geq 0$. Now, if we consider the lower limit $v_{\parallel}(r_{0})=0$, and we combine Equation (12) and Equation (13), we get
@@
with $\psi(r_{0})=0$. In our model, we further assume for simplicity that the magnetic field is purely radial and that $B(r)\propto 1/r^{2}$. This approximation is justified because Pierrard et al. (2001a) demonstrated that incorporating a spiral magnetic field into the kinetic exospheric model alters the radial dependence of the magnetic field without significantly affecting the resulting acceleration. By introducing the factor $\eta\equiv(r_{0}/r)^{2}$, we can therefore rewrite
@@
Figure 1: Comparison of the standard exospheric model’s computed VDFs for electrons at different radial distances in the left panels (a, c, and e) and for protons at the same radial distances in the right panels (b, d, and f) where the $x$-axis is the parallel velocity and the $y$-axis is the perpendicular velocity with respect to the magnetic field direction. The gray areas correspond to empty regions of velocity space.
Let us now assume that the potential difference $\psi(r)$ (assuming that $\psi(r_{0})=0$) is increasing at first and then reaches a maximum at $r_{m,i}$ where the index $i$ corresponds to $e$ or $p$ for electrons or protons, respectively (this index will also be omitted). This situation would imply that some particles cannot reach this maximum because they do not possess enough energy. To distinguish between these particles and those that can instead reach this maximum, we define a velocity threshold characterized by $v_{\parallel}(r_{m})=0$, since any particle with a positive parallel velocity at $r_{m}$ will escape from the potential well. By replacing $r_{0}$ by $r_{m}$ in Equation (12) and Equation (13) and introducing the factor $\mu\equiv(r_{m}/r)^{2}$, we obtain similarly
@@
These equations have been derived by Lemaire & Scherer (1970) for planetary exospheres and are contextualized here in the case of the solar wind in Fig. 1. The VDFs of this figure are reconstructed from the model results obtained using Equation (15) and Equation (16) and the input parameters of Table 1. Panels a and b show the electron and proton VDFs respectively at $r=r_{0}$. Both species are placed in an increasing potential near the Sun because of the mass difference between protons and electrons. Since the electrons are less impacted by the gravity field than the protons, this ultimately results in generating a charge separation electric field (which accelerates the protons and slows down the electrons) to maintain quasi-neutrality and zero net current between protons and electrons. However, for the protons, the dominant effect is still the gravity field near the Sun, placing them in an attractive potential at first. The vertical dashed lines in panels a and b represent the boundary $v_{\parallel}=0$. This implies that the only particles that can have a $v_{\parallel}<0$ at $r=r_{0}$ are the particles that were reflected due to the potential. This is specifically the case for the particles that could not reach $r_{m}$ and had to turn back to the Sun. These particles are represented by the circle dashed line of panel a for electrons or the ellipse dashed line of panel b for protons. Those are directly deduced from Equation (16) with $r_{m,e}\rightarrow\infty$ for electrons and $r_{0}\leq r_{m,p}\leq\infty$ for protons. The gray areas are empty regions of velocity space since we do not consider any pitch angle scattering nor the presence of particles coming from infinity. Panels c and d correspond to the radial distance $r_{0}<r<r_{m,p}$, while panels e and f correspond to the radial distance $r>r_{m,p}$. In panels c and e, the evolution of Equation (15) and Equation (16) for electrons can be seen through the hyperbola and circle dashed lines respectively. The same can be seen for protons in panels d and f, except that here the protons are placed in an attractive potential ($r_{0}<r<r_{m,p}$) and then in a repulsive potential ($r>r_{m,p}$). This is why the ellipse dashed line of panels d transforms into a hyperbola with its foci on the $v_{\parallel}$-axis.
@@
These dashed lines delimit the regions of velocity space labeled with “A”, “B”, and “C” in Fig. 1. The A region corresponds to the particles that do not have sufficient energy to overcome the potential maximum; these are called “ballistic” particles. The B region corresponds to particles that have enough energy to overcome this potential maximum and are therefore called “escaping” particles. The C region corresponds to trapped particles that lack sufficient energy to overcome the potential maximum and cannot reach $r_{0}$. Instead, they are perpetually reflected because of magnetic moment conservation. In principle, such particles cannot originate directly from $r_{0}$. However, as generally assumed (e.g., Maksimovic et al. 1997a; Lamy et al. 2003; Zouganelis et al. 2004), our model still accounts for them because this region is expected to eventually become saturated with particles: any particle entering this region of velocity space (whether through rare collisions or interactions with plasma waves) will remain there indefinitely unless scattered out. Therefore, we consider that the VDF of region C is also described by Equations (6) and (9) for protons and electrons respectively, ensuring VDF continuity between ballistic and trapped particles. Note that in our example there is no region for trapped protons (Fig. 1d), but this is not always the case. In fact, with the right parameters, trapped particles could be obtained close to $r_{0}$ (see, e.g., Lamy et al. 2003).
@@
To compute the VDFs of our model, we must determine the electric potential $V(r)$, since both normalized potentials $\psi_{p}(r)$ and $\psi_{e}(r)$ depend on it. This electric potential is here found self-consistently through the use of the quasi-neutrality between charges and zero-current equations which are assumed to be strictly respected over sufficiently large scales in the solar wind. To derive $V(r)$, we define two key quantities. First, the total normalized potential difference of protons between $r_{m}\equiv r_{m,p}$ and $r_{0}$ is defined as
@@
with $\phi_{m}\equiv\phi(r=r_{m})$, $\phi_{0}\equiv\phi(r=r_{0})$, $V_{m}\equiv V(r=r_{m})$ and $V_{0}\equiv V(r=r_{0})$. Similarly, the normalized potential difference of electrons between $r_{0}$ and $r_{\infty}$ is defined as

### v3

@@ Abstract
Context. Parker Solar Probe (PSP) observations have revealed that most of the solar wind acceleration occurs very close to the Sun. This acceleration is partly due to the global electric potential originating from the mass disparity between electrons and protons, coupled with the constraints of charge quasi-neutrality and zero-current conditions in the solar wind plasma. However, the exact mechanism that accounts for the remaining acceleration has not yet been identified.
Aims. We aim to provide a framework that incorporates the electric-field-driven component of the acceleration while also introducing an additional acceleration mechanism via a velocity-space diffusion of the particles. This will help us determine the extent of extra acceleration, beyond the electric-field-driven component, required to fully reproduce the acceleration of the solar wind in theoretical models.
Methods. We modified an existing kinetic exospheric model to account for the unexplained solar wind acceleration by including velocity-space diffusion, thereby capturing the effect of collisions and wave-particle interactions within the exospheric approach. We compared the electric field derived from the sunward deficit of velocity distribution functions observed by PSP between 13.3 and 50 solar radii ($R_{s}$) with the electric field found self-consistently by the kinetic exospheric model.
Results. The effect of velocity-space diffusion is found to reduce the temperature anisotropy and impact the solar wind acceleration while leaving the electric potential unchanged. The approach described in this work enables the diffusion to be adjusted to effectively reduce or increase the solar wind acceleration. Even without diffusion, the model is able to reproduce the anticorrelation between the electric potential and the solar wind terminal velocity found by PSP. This suggests that the electric potential might still be of major importance in explaining the solar wind acceleration.
@@ 1 Introduction
The electrons and protons that are blown away from the Sun through interplanetary space are the main constituents of a continuous outflow of particles referred to as the solar wind. The exact mechanism that allows the solar wind to reach observed velocities averaging between 300 and 800 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ remains one of the most significant open questions in space physics (e.g., Rouillard et al. 2021). Since the first in situ solar wind observations in 1959, neither magnetohydrodynamic nor kinetic models have been able to self-consistently explain the solar wind acceleration. In particular, magnetohydrodynamic models are not able to capture the origin of the fast solar wind with realistic coronal temperatures and often require additional heating mechanisms (Parker 1965; Leer et al. 1982; Hansteen & Velli 2012). Although some kinetic approaches can provide an explanation for the presence of peculiar observed velocity distribution functions (VDFs) of solar wind particles and propose potential mechanisms of solar wind acceleration (e.g., Hollweg & Isenberg 2002; Vocks et al. 2005; Seough et al. 2015), they generally lack any global electric-field-driven component of the acceleration. We investigated this acceleration by improving a specific class of kinetic model known as exospheric models, which are fundamentally based on considering the electric-field-driven component as the primary acceleration mechanism. A comprehensive review of these models has been provided by Fahr & Shizgal (1983).
@@
The first exospheric models were developed for planetary exospheres and were later adapted for the solar wind (Chamberlain 1960). These models describe the radial evolution of the particle VDFs by neglecting collisions between particles. The first model of the solar wind that included a consistent “ambipolar” electric field was developed by Lemaire & Scherer (1971) for a collisionless regime above the solar surface at a given altitude called the “exobase,” which was chosen to be at 6 solar radii ($R_{s}$). This electric field is considered to be ambipolar in the sense that it satisfies the quasi-neutrality and zero-current conditions between protons and electrons observed to be preserved in the solar wind (e.g., Feldman et al. 1975). The challenging aspect of these models was to determine the radial evolution of this ambipolar electric field self-consistently with the exospheric approach, especially in the case of non-monotonic proton potential (typically found when assuming an exobase below 6 $R_{s}$, which better describes the solar wind case), as noted by Jockers (1970). Lamy et al. (2003), inspired by Jockers (1970), developed a generalized method for approximating the solution for this self-consistent electric potential.
@@
The use of Kappa (also known as Lorentzian) VDFs for electrons, in contrast to Maxwellian VDFs for protons, constitutes one of the major breakthroughs that was first introduced in an exospheric model by Pierrard & Lemaire (1996) and applied to the solar wind by Maksimovic et al. (1997a). This approach was motivated by the need to account for the so-called suprathermal electrons observed extensively in the solar wind (e.g., Lin 1980; Pilipp et al. 1987; Pierrard et al. 2016; Abraham et al. 2022; Salem et al. 2023). Such suprathermal electron populations are better described by a Kappa distribution function rather than a single Maxwellian distribution, as shown, for instance, by Maksimovic et al. (1997b) and more recently by Zheng et al. (2024). These electrons are high-energy particles that do not follow the classical single Maxwellian distribution fitted to the lower-energy part (called the core) of the observed distribution. These suprathermal distributions are typical for the halo, which extends in all directions, and the electron strahl, a more concentrated component observed in the anti-sunward direction (e.g., Feldman et al. 1975; Pilipp et al. 1987; Pierrard et al. 2001b). The mathematical description of Kappa distributions required for exospheric models was first introduced by Pierrard & Lemaire (1996) and generalized for non-monotonic potentials by Lamy et al. (2003). Within the scope of this model, using a Kappa distribution for electrons increases the number of electrons with high speeds in the anti-sunward direction, which effectively increases the electric potential difference between any two radial distances (see Pierrard & Lazar 2010 for a review). Therefore, this allows more protons to reach the maximum of the non-monotonic proton potential (usually within 1 to 5 $R_{s}$), which in turn accelerates the wind to realistically high bulk speeds without requiring unrealistically high temperatures. Following the same approach, Zouganelis et al. (2004) showed that similar results could also be obtained with a sum of two Maxwellians or a sum of a Maxwellian and a Kappa distribution (for the core and strahl), which would not limit the significance of the strahl to the power law factor, $\kappa$.
@@
Given the simplicity and effectiveness of the Lamy et al. (2003) approach in assigning at least part of the unexplained acceleration to the presence of suprathermal electrons, the model was later extended to minor ions (ions with lower abundances than protons) by Pierrard & Pieters (2014). More recently, an improved version of this exospheric model, incorporating regularized Kappa distribution functions (which allow even lower $\kappa$ values and therefore a higher fraction of suprathermal particles), was compared with Parker Solar Probe (PSP) and Solar Orbiter observations by Pierrard et al. (2023). Their results demonstrate that this model reproduces the radial evolution of the solar wind bulk velocity but cannot accurately reproduce the observed anisotropy in the solar wind. This discrepancy arises from the intrinsic conservation of energy and magnetic moment, which strictly constrain modeled VDFs in velocity space.
@@
It has already been noted on multiple occasions that the idealized view provided by the exospheric model is not completely accurate. In particular, collisions and plasma waves are expected to be the mechanisms responsible for the scattering of electrons in the solar wind (Pierrard et al. 2001c; Salem et al. 2003; Pierrard et al. 2011; Boldyrev et al. 2020; Micera et al. 2020; Verscharen et al. 2022; Yoon et al. 2024). These effects inevitably populate regions of velocity space that would otherwise be inaccessible and are believed to be the main mechanisms responsible for the presence of the so-called halo population of higher-energy electrons that are not preferentially anti-sunward directed, unlike the strahl (Feldman et al. 1975; Pilipp et al. 1987; Maksimovic et al. 1997b; Štverák et al. 2008, 2009). Therefore, we propose a way to simulate the effect of particle scattering on the moments of the electron VDF that are modeled by introducing a new, “diffused” population of particles.
@@
We aim to improve the existing kinetic exospheric model by first properly describing the exospheric framework applied to the solar wind. Therefore, Sect. 2 provides a review of an existing kinetic exospheric model. Section 3 describes how we incorporated the velocity-space diffusion within the reviewed kinetic exospheric approach. In particular, we investigate how diffused electrons affect the solar wind temperature, bulk velocity, and density in Sect. 4. The relevance of the kinetic exospheric model is addressed in Sect. 5, as it effectively explains one of the components responsible for solar wind acceleration through its self-consistent electric field, derived from the quasi-neutrality and zero-current conditions. Finally, the conclusions and future perspectives are provided in Sect. 6.
@@ 2 Review of the solar wind kinetic exospheric model
The kinetic exospheric model is based on the assumption that the solar wind becomes collisionless above a radial distance $r_{0}$, called the exobase. This is generally defined by considering that the ratio of the plasma density scale height to the particle mean free path (called the Knudsen number), must be equal to one for this assumption to be valid. More precisely, protons and electrons in the solar wind have different reference altitudes at which their respective Knudsen number reaches unity, as electrons typically have a shorter mean free path than protons (Maksimovic et al. 1997a). However, for simplicity, we did not make this distinction and instead considered a simple, sharp transition between collision-dominated and collisionless regions at $r_{0}=1.2\leavevmode\nobreak\ R_{s}$. This exobase level was arbitrarily chosen to emphasize the model’s physics while remaining generally close to the expected position of the actual exobase above the solar surface (Pierrard et al. 2023). We began our mathematical formulation of the model from the general kinetic equation.
@@
is the force applied to each particle with ${g}$ being the gravitational acceleration, $q_{i}$ the charge of the particle, ${E}$ the electric field, and ${B}$ the magnetic field. The right-hand side of Eq. (1) corresponds to particle–particle interactions. Considering the collisionless assumption in the exosphere (above the exobase), we can reduce this equation to the Vlasov equation,
@@
The Liouville theorem asserts that the VDF is constant along the trajectories of the system. This means that any function $f(c_{1},c_{2},...,c_{n})$, where $c_{1},c_{2},...,c_{n}$ are constants of the motion through external forces ${F}$, is a solution to Eq. (3). In the case of the solar wind, the constants of the motion for each particle are the total (kinetic plus potential) energy,
@@
where $N_{p}$ is a factor of normalization that allows us to retrieve the specified proton density at $r_{0}$ by integrating the proton VDF in velocity space, $w_{p}=\sqrt{2k_{b}T_{0p}/m_{p}}$ is the proton thermal velocity with $k_{b}$ being the Boltzmann constant, $T_{0p}$ being the proton temperature at $r_{0}$, $m_{p}$ being the proton mass, and $\mathcal{E}_{p}/(k_{b}T_{0p})$ is the normalized proton energy such that
@@
Here, $\kappa$ is the parameter associated with a power law decreasing Kappa VDF, instead of an exponentially decreasing Maxwellian VDF (i.e., for $\kappa\rightarrow\infty$ the Maxwellian distribution is retrieved); $w_{e}=\sqrt{(2\kappa-3)k_{b}T_{0e}/(\kappa m_{e})}$ is the electron thermal velocity, $T_{0e}$ is the electron temperature at $r_{0}$, $m_{e}$ is the electron mass, $\Gamma$ is the Gamma function, $N_{e}$ is a factor of normalization (defined in Sect. 3) that allows us to retrieve the specified electron density at $r_{0}$ by integrating the electron VDF in velocity space, and $\mathcal{E}_{e}/[(\kappa-3/2)k_{b}T_{0e}]$ is the normalized electron energy such that
@@
being the normalized total potential for electrons. We note that $f_{p}$ and $f_{e}$ are functions of $\mathcal{E}$ but also $\mathcal{M}$ since the magnetic moment will have an influence on the allowed perpendicular velocities of particles depending on their origin.
@@
These equations are given in a general form for electrons and protons, as noted by the species index $i$, which will be omitted, for convenience, in the remainder of the description. If we assume that the reference altitude, $r_{0}$, is exactly the point where the solar wind becomes collisionless and that below $r_{0}$ the particles are dominated by collisions (in such a way that there is no preferential direction for the particle motion), then at $r_{0}$ all particles should be constrained to $v_{\parallel}(r_{0})\geq 0$. Now, if we consider the lower limit $v_{\parallel}(r_{0})=0$, and we combine Eqs. (12) and (13), we get
@@
with $\psi(r_{0})=0$. In our model, we further assume, for simplicity, that the magnetic field is purely radial and that $B(r)\propto 1/r^{2}$. This approximation is justified because Pierrard et al. (2001a) demonstrated that incorporating a spiral magnetic field into the kinetic exospheric model alters the radial dependence of the magnetic field without significantly affecting the resulting acceleration. By introducing the factor $\eta\equiv(r_{0}/r)^{2}$, we can therefore rewrite
@@
Figure 1: Comparison of the standard exospheric model’s computed VDFs for electrons at different radial distances (panels a, c, and e) and for protons at the same radial distances (panels b, d, and f). The $x$-axis is the parallel velocity, and the $y$-axis is the perpendicular velocity with respect to the magnetic field direction. The gray areas correspond to empty regions of velocity space.
Let us now assume that the potential difference $\psi(r)$ (assuming that $\psi(r_{0})=0$) is increasing at first and then reaches a maximum at $r_{m,i}$ where the index $i$ corresponds to $e$ or $p$ for electrons or protons, respectively (this index will also be omitted). This situation would imply that some particles cannot reach this maximum because they do not possess enough energy. To distinguish between these particles and those that can instead reach this maximum, we defined a velocity threshold characterized by $v_{\parallel}(r_{m})=0$, since any particle with a positive parallel velocity at $r_{m}$ will escape from the potential well. By replacing $r_{0}$ with $r_{m}$ in Eqs. (12) and (13) and introducing the factor $\mu\equiv(r_{m}/r)^{2}$, we similarly obtain
@@
These equations were derived by Lemaire & Scherer (1970) for planetary exospheres and are contextualized here in the case of the solar wind in Fig. 1. The VDFs of this figure are reconstructed from the model results obtained using Eqs. (15) and (16) and the input parameters of Table 1. Panels a and b show the electron and proton VDFs, respectively, at $r=r_{0}$. Both species are placed in an increasing potential near the Sun because of the mass difference between protons and electrons. Since the electrons are less impacted by the gravity field than the protons, this ultimately results in generating a charge separation electric field (which accelerates the protons and slows the electrons) to maintain quasi-neutrality and zero net current between protons and electrons. However, for the protons, the dominant effect is still the gravity field near the Sun, which places them in an attractive potential at first. The boundary at $v_{\parallel}=0$ implies that the only particles that can have a $v_{\parallel}<0$ at $r=r_{0}$ are the particles that were reflected due to the potential. This is specifically the case for the particles that could not reach $r_{m}$ and had to turn back to the Sun. These particles are directly deduced from Eq. (16) with $r_{m,e}\rightarrow\infty$ for electrons and $r_{0}\leq r_{m,p}\leq\infty$ for protons. There are empty regions of velocity space since we did not consider any pitch angle scattering nor the presence of particles coming from infinity. Panels c and d correspond to the radial distance $r_{0}<r<r_{m,p}$, while panels e and f correspond to the radial distance $r>r_{m,p}$. In panels c and e, the evolution of Eqs. (15) and (16) for electrons is represented by dashed lines forming hyperbolas and circles, respectively. The same can be seen for protons in panels d and f, except that here the protons are placed in an attractive potential ($r_{0}<r<r_{m,p}$) and then in a repulsive potential ($r>r_{m,p}$). This is why the ellipse dashed line of panel d transforms into a hyperbola with its foci on the $v_{\parallel}$-axis.
@@
These dashed lines delimit the regions of velocity space labeled A, B, and C in Fig. 1. The A region corresponds to the particles that do not have sufficient energy to overcome the potential maximum; these are called “ballistic” particles. The B region corresponds to particles that have enough energy to overcome this potential maximum and are therefore called “escaping” particles. The C region corresponds to trapped particles that lack sufficient energy to overcome the potential maximum and cannot reach $r_{0}$. Instead, they are perpetually reflected because of magnetic moment conservation. In principle, such particles cannot originate directly from $r_{0}$. However, as generally assumed (e.g., Maksimovic et al. 1997a; Lamy et al. 2003; Zouganelis et al. 2004), our model still accounts for them because this region is expected to eventually become saturated with particles: any particle entering this region of velocity space (whether through rare collisions or interactions with plasma waves) will remain there indefinitely unless scattered out. Therefore, we consider that the VDF of region C is also described by Eqs. (6) and (9) for protons and electrons, respectively, ensuring VDF continuity between ballistic and trapped particles. We note that, in our example, there is no region corresponding to trapped protons (Fig. 1d); however, this is not always the case. In fact, with the right parameters, trapped particles could be obtained close to $r_{0}$ (see, e.g., Lamy et al. 2003).
@@
To compute the VDFs of our model, we had to determine the electric potential $V(r)$ since both the normalized potentials, $\psi_{p}(r)$ and $\psi_{e}(r),$ depend on it. Here, this electric potential is found self-consistently through the use of the quasi-neutrality between charges and zero-current equations, which are assumed to be strictly respected over sufficiently large scales in the solar wind. To derive $V(r)$, we defined two key quantities. First, the total normalized potential difference of protons between $r_{m}\equiv r_{m,p}$ and $r_{0}$ was defined as
@@
with $\phi_{m}\equiv\phi(r=r_{m})$, $\phi_{0}\equiv\phi(r=r_{0})$, $V_{m}\equiv V(r=r_{m})$ and $V_{0}\equiv V(r=r_{0})$. Similarly, the normalized potential difference of electrons between $r_{0}$ and $r_{\infty}$ was defined as